# From Canopy to Ground via ForestGen3D: Learning Cross-Domain Generation of 3D Forest Structure from Aerial-to-Terrestrial LiDAR

### v1

@@ 3.4 Plot-scale ALS-Guided Generation
Building on these qualitative comparisons, we also present Figure 11, which illustrates biophysical tree metrics derived from ALS, TLS, ALS+TLS and ALS+ForestGen3D data sources. This figure offers a histogram per metric (one per sub-column) depicting the distribution of key biophysical tree attributes, offering insight into the variability of tree structures within the sampled region. Such histograms were computed over the individual plots of 25 meter radius. To produce these histograms, we evaluate tree height, diameter at breast height (DBH), crown diameter (CrD), and crown volume (CrV) biometrics using established deep learning-based segmentation methods [19]. These segmentation models first segment each point cloud inside the detected vegetation 3D bounding box, into its structural components: ground, low vegetation, stems, branches and foliage, which are then used as input for geometric fitting techniques that yield the final biophysical metrics.
Tree height was estimated using the highest elevation point in a detected tree and subtracting from it a projection of that point into the ground surface. To estimate DBH, we used the method of [24] on the segmented stem, modifying it for robustness with a RANSAC based linear interpolation of fitted circles along multiple heights around breast height. Note that the RANSAC based fit was used to mitigate the effects of outliers. Crown diameter was estimated by projecting the segmented foliage points onto a 2D plane (collapsing the elevation/altitude dimension) and measuring the maximally distant points in the projection. For crown volume, we employ a $k$-means-based convex hull approach, as described in [25]. This method partitions the segmented foliage points into $k$-clusters of closely packed 3D points, computes the convex hull of each cluster, and sums the volumes of the $k$-clusters. This approach mitigates the effect of including empty-space volume in the calculation, which can occur when applying a single convex hull computation to the full crown structure.
Figure 11 highlights the complementary strengths and inherent limitations of different laser scanning techniques. Among these, the integration of ALS and TLS in the third row across all three examples, offers the most comprehensive view of forest structure, leveraging the wide-area coverage of ALS with the high-resolution, near-ground detail of TLS. This combination enables high-fidelity reconstruction across the entire vertical profile of forest stands above, within, and below the canopy. As such, metrics derived from this data are considered the gold standard reference for evaluating performance across all key biophysical metrics. Individually, ALS in the first row provides efficient and scalable coverage of large landscapes, capturing canopy-top structure effectively from an overhead perspective resulting in estimates without missing information issues for tree height and crown diameter. However, relative to TLS or the ALS+TLS benchmark, ALS alone lacks the resolution necessary to characterize lower-canopy and understory features due to occlusion. This limitation results in poor performance for DBH, with a highly sparse distribution across almost all example cases. Height and crown volume metrics are relatively close to those from the ALS+TLS distributions as they do not heavily depend on the understory information. TLS in the second row across all examples offers fine-grained structural information along the tree-stand direction which results in a denser DBH distribution compared to ALS. In terms of crown diameter we see that the range values in TLS are generally contained within those ALS-derived, this corroborates at the crown level, our assumption that TLS crown points are contained within the corresponding crown points from ALS. In the case of height, the distribution of the TLS-derived height is also generally contained or similar in range relative to the corresponding ALS-derived. A few heights in the smaller range appear in TLS compared to ALS, coming likely from vegetation/trees under taller trees that ALS did not capture due to occlusion. For the crown volume, a comparison between the ALS and TLS distribution varies since ALS is good at capturing the crown in the lat/long direction whereas TLS is known to capture the crown better in the vertical/elevation direction. Our integrated ALS + ForestGen3D approach in the last row, built with the objective of compensating the structural blind spots of ALS bridges the gap between airborne-scale reach and terrestrial-scale resolution. As observed in the representative examples, the improvements provided by our model produces distributions of biophysical metrics that better approximates the distributions of the gold-standard ALS+TLS. Note also, that in general, the tree height and crown diameter remains approximately contained within that of the corresponding ALS-derived distributions supporting our containment guarantees under ELBO.
(a) Ex.1 Height (col1), DBH (col2), CrD (col3) and CrV (col4) distribution
(b) Ex.2 Height (col1), DBH (col2), CrD (col3) and CrV (col4) distribution
(c) Ex.3 Height (col1), DBH (col2), CrD (col3) and CrV (col4) distribution
(d) Ex.4 Height (col1), DBH (col2), CrD (col3) and CrV (col4) distribution
Figure 11: Biophysical metric distributions for tree height, DBH, crown diameter and crown volume including comparisons between ALS (row1), TLS (row2), ALS+TLS (row3) and ALS+ForestGen3D (row4).
Distributional differences in key biophysical metrics: tree height, DBH, CrD, and CrV were also quantified to evaluate model performance. Given the sparse nature of plot-scale biophysical metric distributions (typically 10–30 trees per plot), we employed the Wasserstein distance (WD), which is particularly well suited for such data. This metric offers advantages over alternatives like the Kullback–Leibler divergence, which can become unstable when there is limited overlap between distributions. The Wasserstein distance provides an intuitive and robust measure of distance between two distributions by quantifying the minimum effort required to transform one into the other, and is often also referred to as the Earth Mover’s Distance. We make the distinction, and refer to the metric measuring distance between biometric distributions and the one computing distance over point-clouds for reconstruction error as the EMD. Table 2, compares WD distances between ALS, TLS, and ALS+TLS distributions relative to ALS+ForestGen3D. Each example corresponds to the distributions shown in Figures 11(a), 11(c), and 11(d). Overall, these results consistently show the smallest Wasserstein distances between ALS+ForestGen3D and ALS+TLS, highlighting the model’s effectiveness in generating realistic below-canopy structural information. This leads to biophysical metric distributions that closely resemble those derived from the more comprehensive ALS+TLS data. Further, the distances between ALS+ForestGen3D and the standalone ALS or TLS sources reveal a consistent trend: smaller distances are observed for the source that provides greater information content for the 3D spatial features associated with a given metric. For instance, the distance for the height metric is smaller between ALS+ForestGen3D and ALS than between ALS+ForestGen3D and TLS. This is expected, as ALS is less affected by occlusion in both the ground and crown top regions, making it more uniform and informative for height-related measurements. Finally, note that the scale of the CrV metric is significantly higher than those of height, DBH, and CrD. This is due to the fact that CrV is measured in cubic meters, whereas the other metrics are measures expressed in meters.
Table 2: Plot-scale sensing source biometric distribution differences relative to ALS + ForestGen3D.
@@
Example 1 Example 2 Example 3
Sensing Source Height DBH CrD CrV Height DBH CrD CrV Height DBH CrD CrV
ALS 1.04 0.30 0.79 117.94 0.25 0.18 0.49 44.91 0.29 0.24 1.14 24.20
TLS 1.94 0.23 1.25 98.93 0.30 0.23 0.75 24.82 0.31 0.11 0.84 48.18
ALS + TLS 0.80 0.21 0.85 80.23 0.17 0.18 0.52 11.58 0.12 0.15 0.32 26.8
To generalize these findings, Table 3 summarizes the average Wasserstein distances computed over the 51 plots in the testset. This expanded analysis reinforces the trends observed in the examples, confirming that ALS+ForestGen3D distributions most closely match those from ALS+TLS across key biometrics. The consistency of these results across a large number of diverse plots further supports the overall good performance capacity of the ForestGen3D model when the characteristics of the mixed conifer ecosystem remain more or less similar.
Table 3: Overall biometric distribution differences between sensing sources relative to ALS + ForestGen3D.
@@
Sensing Source Height DBH CD CV
ALS 1.04 0.30 0.79 117.94
TLS 1.94 0.23 1.25 98.93
ALS + TLS 0.80 0.21 0.85 80.23
@@ 3.5 Regional-scale ALS-Guided Generation
To evaluate the scalability of the ForestGen3D model, we applied it to generate 3D forest structure over a larger spatial extent, a 200-meter radius region centered at a location where a TLS scan is available. While the TLS scan itself covers a range of only  25 meters, this broader area enables us to test the model’s ability to extrapolate beyond small-scale. Notably, the ALS data used in this region was not part of the training set, providing a rigorous test of the model’s generalization capabilities. The generation procedure at this larger scale mirrors the approach used for plot-level synthesis. First, we apply a tree detection algorithm across the entire ALS region of interest to identify areas with occlusion when viewed from above, which are typically regions where airborne LiDAR fails to resolve understory structure. Second, we extract 3D bounding boxes around these detected trees and apply the ForestGen3D model independently within each bounding box using the local ALS point cloud as input. Finally, the resulting synthetic point clouds are aggregated over the full 200-meter radius area keeping track of their localization to produce a coherent forest-scale 3D reconstruction. Figure 12 illustrate representative snapshot examples of the resulting generation/reconstruction. In these visualizations, point clouds are color-coded as follows: magenta for ALS, turquoise for TLS, peach for ALS + ForestGen3D generation, and white for the generated ForestGen3D sample alone. In this test case where TLS data is not available, the model generates understory structure consistent with the characteristic statistical distribution observed in the training data.
@@
Importantly, these large-scale results reflect and reinforce the findings from the plot-scale experiments. ForestGen3D generates foliage and branching structures that blend smoothly into those observed in the ALS data, maintaining spatial consistency and fidelity. In understory regions completely lacking ALS points, the model synthesizes realistic and plausible tree trunks in locations and with diameters that appear structurally consistent with the associated crown geometry provided by ALS. When sparse ALS points are present in the understory, the generated output enhances and complements this limited data by reinforcing plausible stem/branch geometry. At the ground level, the model enriches the representation of shrubs and low vegetation while also improving the resolution of ground returns and the near-ground structure.
@@
Figure 13: Distribution of distances from generated points to the ALS convex envelope across 10 landscape-scale plots. Each violin plot corresponds to a 200-meter radius region labeled along the x-axis. The y-axis indicates the minimum Euclidean distance from generated points lying outside the ALS convex hull to their nearest point on the hull surface. The violin shape represents the distribution of these distances; the internal black bar denotes the median. Annotations above each plot report: NN, the number of trees generated in the region; Out, the fraction of generated points falling outside the ALS envelope; and the mean ± standard deviation ($\mu\pm\sigma$) of out-of-envelope distances. Across all regions, the proportion of outliers remains below 1.6%, and deviations remain small, typically under 0.2m, indicating strong spatial consistency between the generated structure and the ALS-derived envelope, thus supporting our EPC under ELBO remark.
@@
To quantitatively assess the generation performance of our model at this landscape scale, we analyzed the EPC metric and the distribution of distances between generated points falling outside the ALS convex envelope and their nearest point on the envelope surface. Figure 13 summarizes this analysis across 10 evaluation plots, each representing a 200 meter radius ALS-only region labeled by plot name on the x-axis. Each violin plot shows the distribution of out-of-envelope distances, computed only for generated points that lie outside the ALS convex hull. The line inside each violin represents the median, while the shape reflects the underlying distance histogram. Text annotations above each violin indicate: (i) $N$, the number of trees generated per 200m radius region; (ii) Out, the ratio of generated points outside the ALS convex hull (relative to the total number of points); and (iii) the mean ± standard deviation of the out-of-envelope distances. Across all plots, the fraction of points outside the ALS envelope remains consistently low, ranging from 1.4% to 1.6%, and mean distances remain below 0.2m. This confirms that the generative model is contained within the implicit ALS spatial convex hull constraints, even in landscape-scale scenarios demonstrating good quality generations via the EPC. The narrow shapes and small deviations provide strong empirical support for our containment assumption and demonstrate that the DDPM model preserves structural plausibility across diverse spatial contexts.

### v2

@@ 3.5.1 Do ForestGen3D-Derived Biometrics Approximate TLS-Derived Distributions?
In the first set of experiments for this task, we compute biophysical tree metric distributions derived from ALS, TLS, ALS+TLS and ALS+ForestGen3D data sources. For this, we use a histogram per metric (one per column) depicting the distribution of key biophysical tree attributes, offering insight into the variability of tree structures within the sampled plot region. Such histograms were computed over co-registered sources at individual plots of 25 meter radius. To produce these histograms, we evaluate tree height, diameter at breast height (DBH), crown diameter (CrD), and crown volume (CrV) using established deep learning–based tree segmentation methods [26]. These models segment each point cloud within detected vegetation bounding boxes into structural components: ground, low vegetation, stems, branches, and foliage, which are then used as inputs to geometric fitting procedures that yield the final biophysical metrics. Tree height is estimated using the highest elevation point within each detected tree, subtracting a projection of that point onto the ground surface. DBH is computed following the method of [34] applied to segmented stems, with robustness improved through a RANSAC based linear interpolation of fitted circles across multiple heights around breast height to mitigate the effect of outliers. Crown diameter is estimated by projecting segmented foliage points onto a horizontal plane and measuring the maximum pairwise distance. Crown volume is computed using a $k$-means–based convex hull approach [35], which partitions foliage points into spatial clusters, computes the convex hull of each cluster, and sums their volumes. This approach mitigates overestimation caused by empty space inclusion when applying a single convex hull to the full crown.
In Figure 11, integrated ALS+TLS data (third row across all examples) provides the most comprehensive representation of forest structure by combining ALS’s canopy coverage from the top with TLS’s detailed near-ground sampling. As such, biometrics derived from ALS+TLS are treated as a gold-standard reference for comparison. ALS-only data (first row) efficiently captures canopy-top structure and yields reliable estimates of tree height and crown diameter, but lacks the resolution necessary to characterize lower-canopy and understory features due to occlusion. This limitation leads to sparse and biased DBH estimates across most examples, while height remains closer to the ALS+TLS reference because it depend less directly on stem-level information. TLS-only data (second row) provides dense sampling of stems and lower-canopy elements, resulting in well-defined DBH distributions. TLS-derived crown diameter and height distributions are generally contained within ALS-derived ranges, consistent with the assumption that TLS crown points are spatially embedded within the ALS envelope. Smaller height values observed in TLS reflect understory trees occluded in ALS data or missed upper canopy. Differences in crown volume distributions reflect the complementary directional sampling strengths of ALS (horizontal canopy extent) and TLS (vertical crown structure). The ALS+ForestGen3D results (last row) indicate that the proposed model compensates for ALS structural blind spots while preserving ALS derived landscape geometry. The resulting biometric distributions consistently approximate those of the ALS+TLS reference across all evaluated metrics. Improvements are most pronounced for DBH and crown volume, indicating recovery of sub-canopy and stem structure absent in ALS-only data. Importantly, tree height and crown diameter distributions remain largely contained within the corresponding ALS-derived ranges, supporting the model’s geometric containment guarantees under the ELBO objective. Together, these results indicate that ForestGen3D generates structurally plausible 3D vegetation consistent with TLS-derived distributions while maintaining fidelity to ALS-provided canopy geometry. Other representative example illustrations can be found in Figures 15, 16 and 17 in the Appendix.
Figure 11: Plot-scale distributions of biophysical tree metrics derived from different sensing sources. Rows correspond to ALS (row 1), TLS (row 2), ALS+TLS (row 3), and ALS+ForestGen3D (row 4). Columns show distributions for tree height (col. 1), diameter at breast height (DBH; col. 2), crown diameter (CrD; col. 3), and crown volume (CrV; col. 4). Each histogram represents metrics computed over a 25 m radius plot. ALS+TLS is treated as the structural reference, while ALS+ForestGen3D closely approximates its metric distributions, indicating recovery of sub-canopy structure while preserving ALS-derived spatial constraints.
Beyond qualitative assessment, we quantitatively evaluate model performance in this downstream task, by comparing the distributions of tree height, diameter at breast height, crown diameter, and crown volume derived from the different 3D sources. Because plot-scale biometric distributions are typically sparse (on the order of 10–30 trees per plot), we adopt the Wasserstein distance (WD) as our primary distributional comparison metric. WD is particularly well suited for this domain, as it remains stable under limited sample sizes and partial support overlap, unlike divergence-based measures such as the Kullback–Leibler divergence. The Wasserstein distance (see [36] for an overview) provides an intuitive measure of dissimilarity by quantifying the minimum cost required to transform one distribution into another and is therefore also referred to as the Earth Mover’s Distance. We explicitly distinguish this use of WD for comparing biometric distributions from the EMD metric which operates directly on 3D point sets.
Table 2: Average Wasserstein distances between plot-scale biophysical metric distributions derived from ALS, TLS, and ALS+ForestGen3D relative to ALS+TLS, computed over all 51 plots in the test set. Results confirm that ALS-ForestGen3D yields biometric distributions that most closely approximate those derived from ALS+TLS across all evaluated metrics, demonstrating consistent performance across diverse plots within the mixed-conifer ecosystem.
@@
Sensing Source Height DBH CrD CrV
ALS 0.94 0.29 0.81 $4.87^{3}$
TLS 1.85 0.22 1.11 $4.33^{3}$
ALS + ForestGen3D 0.82 0.205 0.88 $4.24^{3}$
Table 2 summarizes average Wasserstein distances computed across all 51 plots in the test set. Across all examples and metrics, ALS+ForestGen3D consistently exhibits the smallest Wasserstein distances relative to ALS+TLS. Note in general, that in terms of distances the smallests’ are observed relative to the sensing source that provides greater information content for the structural features underlying that metric. For example, Wasserstein distances for tree height are generally smaller between ALS+TLS and ALS than between ALS+TLS and TLS, reflecting the fact that ALS offers more uniform and less occlusion affected sampling of canopy tops and ground surfaces. Conversely, metrics such as DBH, which depend strongly on stem level sampling, show closer agreement between ALS+TLS and TLS. Overall, the results here indicates that the biometric distributions derived from ALS+ForestGen3D generation most closely approximate those obtained from the combined ALS+TLS data, which provides the most complete characterization of forest structure. These findings support that ForestGen3D effectively recovers below-canopy structural information missing from ALS-only observations while remaining consistent with ALS-provided canopy geometry. Note that WD for crown volume is shown in cubic meters, hence its high value difference compared to the other metrics.
The second portion of the biometric estimation task directly assesses the added value of ForestGen3D relative to conventional ALS-based allometric approaches. This consists in estimating DBH from ALS-derived crown diameter and tree height using established allometric relationships following Jucker et al. [37]. Model parameters were trained using 453 matched trees from the Ft. Stewart 2022 dataset, drawn from geographically distinct plots relative to the test set but within the same landscape and vegetation regime. Evaluation was performed on 246 test matched trees from the Ft. Stewart 2024 dataset, with ALS+TLS point clouds used as reference. To mitigate nonuniform sampling effects associated with fixed single scan TLS acquisitions, each point cloud was truncated to a 25 m radius around the TLS scan center. Trees were matched across co-registered ALS, TLS, ALS+TLS, and ALS+ForestGen3D sources using a 0.5 m spatial neighborhood. Table 3 summarizing RMSE’s across all four evaluated biometrics: tree height, DBH, crown diameter, and crown volume show that ALS+ForestGen3D achieved the lowest RMSE. In particular, DBH error was reduced from 0.208 m (ALS allometry [37]) to 0.112 m (ALS+ForestGen3D), while crown volume error was reduced by more than 50%. These results demonstrate that ForestGen3D not only reconstructs realistic 3D structure but also improves downstream biometric estimation relative to conventional ALS based allometric models.
Table 3: Root mean square error (RMSE) of biophysical metrics derived from different sensing sources relative to the ALS+TLS reference. Metrics include tree height, diameter at breast height (DBH), crown diameter (CrD), and crown volume (CrV). DBH estimates for ALS-only inputs are obtained using an allometric model based on crown diameter and height following [37]. Results are reported for a test of 265 matched trees across all sources within the Ft. Stewart 2024 dataset. Lower RMSE values indicate closer agreement with the ALS+TLS benchmark, with ALS+ForestGen3D achieving the lowest error across all metrics.
@@
Sensing Source Height DBH CrD CrV
ALS 0.531 0.208 (allometric [37]) 0.566 $5.112^{3}$
TLS 1.508 0.143 0.6127 $4.029^{3}$
ALS + ForestGen3D 0.480 0.112 0.561 $2.109^{3}$
@@ 3.6 Regional-scale ALS-Guided Generation
To evaluate the scalability of the ForestGen3D model, we applied it to generate 3D forest structure over a larger spatial extent, a 200-meter radius region centered at TLS scan geo-referenced locations. While the TLS scan itself covers a range of only  25 meters, this broader area enables us to test the model’s ability to extrapolate beyond small-scale. Notably, the ALS data used in this region was not part of the training set. The generation procedure at this larger scale mirrors the approach used for plot-level synthesis. First, we apply a tree detection algorithm across the entire ALS region of interest to identify areas with occlusion when viewed from above, which are typically regions where airborne LiDAR fails to resolve understory structure. Second, we extract 3D bounding boxes around these detected trees and apply the ForestGen3D model independently within each bounding box using the local ALS point cloud as input. Finally, the resulting generated point clouds are aggregated over the full 200-meter radius area keeping track of their localization to produce a coherent forest 3D reconstruction. Figure 12 illustrate representative snapshot examples of the resulting generation/reconstruction. In these visualizations, point clouds are color-coded as follows: magenta for ALS, turquoise for TLS, peach for ALS + ForestGen3D generation, and white for the generated ForestGen3D samples alone. In this test case where TLS data is not available, the model generates understory structure consistent with the characteristic statistical distribution observed in the training data.
@@
Importantly, these large-scale results corroborate the findings from the plot-scale experiments. ForestGen3D generates foliage and branching structures that blend smoothly into those observed in the ALS data, maintaining spatial consistency and fidelity. In understory regions lacking ALS points, the model synthesizes realistic and plausible tree trunks in locations and with diameters that appear structurally consistent with the associated crown geometry provided by ALS. When sparse ALS points are present in the understory, the generated output enhances and complements this limited data by reinforcing plausible stem/branch geometry. At the ground level, the model enriches the representation of shrubs and low vegetation while also improving the resolution of ground returns and the near-ground structure.
To explicitly evaluate whether the generated 3D structure remains spatially constrained by the ALS observations, particularly in the absence of TLS ground truth at landscape scales, we next introduce a complementary evaluation based on the expected point containment (EPC) in Eq. (1). This analysis quantifies the extent to which generated points respect the implicit spatial bounds defined by the ALS convex hull, providing a geometry aware validation of large scale structural plausibility.
@@
Figure 13: Distribution of point-to-envelope distances between ForestGen3D generated points and the ALS derived convex hull across ten landscape-scale (200 m radius) regions. The x-axis represents diferent geo-located areas, the y-axis indicates the minimum Euclidean distance from generated points lying outside the ALS convex hull to their nearest point on the hull surface. The violin shape represents the distribution of these distances; the internal black bar denotes the median. Annotations above each plot report: NN, the number of trees generated in the region; Out, the fraction of generated points falling outside the ALS envelope; and the mean ± standard deviation ($\mu\pm\sigma$) of out-of-envelope distances. Across all regions, the proportion of outside points remains below 1.6%, and deviations remain small (typically <0.2 m). These results show that ForestGen3D generates structure that remains spatially consistent with the ALS envelope while enriching sub-canopy detail, confirming that the model respects the geometry encoded in ALS measurements.
@@
For this, we analyzed the EPC metric and the distribution of distances between generated points falling outside the ALS convex envelope and their nearest point on the envelope surface. Figure 13 summarizes this analysis across 10 evaluation plots, each representing a 200 meter radius ALS-only region labeled by plot name on the x-axis. Each violin plot shows the distribution of out-of-envelope distances, computed only for generated points that lie outside the ALS convex hull. The line inside each violin represents the median, while the shape reflects the underlying distance histogram. Text annotations above each violin indicate: (i) $N$, the number of trees generated per 200m radius region; (ii) Out, the ratio of generated points outside the ALS convex hull (relative to the total number of points); and (iii) the mean ± standard deviation of the out-of-envelope distances. Across all plots, the fraction of points outside the ALS envelope remains consistently low, ranging from 1.4% to 1.6%, and mean distances remain below 0.2m, demonstrating good quality generations via the EPC. The results in Fig. 13 provide an important consistency check linking landscape scale ALS geometry with the fine scale structure generated by ForestGen3D. Because ALS often undersamples subcanopy regions, a generative model could in principle hallucinate structure that violates spatial constraints or extends far beyond the areas supported by ALS evidence. The violin plot analysis quantitatively shows that this does not occur in practice: the model generates new subcanopy detail while remaining tightly confined to the ALS spatial envelope. This demonstrates that ForestGen3D enriches incomplete ALS point clouds without distorting broader landscape geometry and, consistent with earlier findings linking high EPC to low Chamfer Distance (CD) and Earth Mover’s Distance (EMD) under TLS supervision, indicates that the model achieves low 3D reconstruction error and high-quality generation at the landscape scale.
Summarizing, across all evaluations, ForestGen3D demonstrates strong and consistent quantitative performance in reconstructing sub-canopy forest structure from ALS input alone. At the individual tree scale, the model achieves low reconstruction error relative to TLS references, as shown by Chamfer distance, Earth’s Moving Distance, and expected point containment (Fig. 7, Fig. 8, Table 1). Biophysical attributes derived from generated point clouds tree height, DBH, crown diameter, and crown volume closely track TLS measurements and outperform or complement ALS-based or ALS-allometry based estimates (Table 2, Table 3), indicating that the model recovers realistic structural detail not observable from ALS alone. At landscape scale, ForestGen3D preserves global spatial consistency with ALS geometry, with the EPC metric and out-of-envelope distance distributions (Fig. 13) confirming that more than 98% of generated points remain contained within the ALS convex hull and deviations remain minimal (<0.2 m). Together, these quantitative analyses demonstrate that ForestGen3D provides a reliable and ecologically coherent enhancement of ALS-only forest structure, producing high-fidelity 3D reconstructions across individual-tree, plot-scale and landscape-level settings.